A study about the number of sleeping hours among call center agents is being conducted in abc company. The population shows a normal distribution with a mean of 6.2 hours. Twentt-five employees were subjected to test and results showed a standard deviation of 0.8. What is the probability that the agents sleep less than 6.5 hours?

1
Expert's answer
2022-04-21T06:34:04-0400

"P(X<6.5)=P(Z<\frac{6.5-6.2}{0.8/\sqrt{25}})=P(Z<1.88)=0.9699"